OpenStudy (anonymous):

Two oppositely charged particles are released from a distance L away from each other, at what time t will they meet? Use conservation Laws. Any help, greatly appreciated.

OpenStudy (anonymous):

Coulomb's Laws

OpenStudy (anonymous):

what's their masses ?

OpenStudy (anonymous):

mass of electron = 9.1e-31kg, and proton = 1.673e-31kg

OpenStudy (anonymous):

so they are the electron and proton? note that when they collide, their potential energies are zero. you can use the following formula \(\Sigma W=E_k\) , since u=0 \(\frac{1}{2} mv^2= \int_0^R -\frac{kQ_1Q_2}{r^2} dr\) for this. and also the linear momentum velcity eq, \(v_1=\frac{m_1-m_2}{m_1+m_2} u_1 +\frac{2m_2}{m_1+m_2} u_2\) for this.

OpenStudy (anonymous):

sorry, the last one won't work. You'll have to use \(m_1 v_1 = -m_2 v_2\) by conservation of momentum

OpenStudy (anonymous):

The mass of the proton provided above is incorrect. M proton =1.672xE-27 Kg